what is 12.5% of 96?
60 Is 75% of what number?
6 Is what percent of 5?

Answer by stanbon(75887)

what is 12.5% of 96?
0.125*x = 96
x = 758
----------------------
60 Is 75% of what number?
60 = 0.75x
x = 80
-----------------------
6 Is what percent of 5?
6 = x*5
x = 6/5 = 1.20 = 120%
=========================
Cheers,
Stan H.